OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN. which of the following graphs represents the equation y-2=3(x + 1) ?

OpenStudy (anonymous):

@Chelsea04 @jim_thompson5910 @Kuoministers @UnkleRhaukus

OpenStudy (anonymous):

I got D but idk

jimthompson5910 (jim_thompson5910):

solve for y y-2=3(x + 1) y-2=3x + 3 y = 3x + 3 + 2 y = 3x + 5 So this graph will have a y-intercept of 5 and a slope of 3

jimthompson5910 (jim_thompson5910):

oh it's -1

jimthompson5910 (jim_thompson5910):

y-2=3(x - 1) y-2=3x - 3 y = 3x - 3 + 2 y = 3x - 1 So this graph will have a y-intercept of -1 and a slope of 3

OpenStudy (anonymous):

so is it d?

jimthompson5910 (jim_thompson5910):

in graph D, the y-intercept is -4 or the point (0,-4), so that's not correct

OpenStudy (anonymous):

how do u know which onme it is @Chelsea04

jimthompson5910 (jim_thompson5910):

which graph has a y-intercept of (0,-1) ?

OpenStudy (anonymous):

b! :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

very good